The analysis of a hydrocarbon revealed that it was 85.6281% C and 14.3719% H by mass. When 2.94 g of the gas was stored in a 1.4 L flask at −32.1762C, it exerted a pressure of 450 Torr. What is the molecular formula of the hydrocarbon?
1
Expert's answer
2012-03-30T07:52:42-0400
Mol C = 85.6281 g / 12.011 = 7.1291
Moles H = 14.3719 g / 1.00794 = 14.259
we divide by the smallest number empirical formula CH2 (mass = 14.0269 g/mol)
p = 450/760 = 0.592 atm
T = −32.1762 + 273 = 240.8238 K
moles = pV / RT = 0.592 *1.4/ 0.0831* 240,8238 = 0.0414
Molar mass = 2.94 g / 0.0414 = 71.014 g/mol
71.014 / 14.0269 = 5
we multiply the empirical formula by 5 and we get C5H10